vinyl disc is 12 inches in diameter. when a record player needle is placed on the outside edge of a vinyl record, usa a calculator to find how far the needle travels in one rotation. Round to the nearest tenth.
Answer by checkley71(8403)

CIRCUMFERANCE OF A CIRCLE=2*PI*R=2*3.14*6=37.68 INCHES IS THE ANSWER.